# nsarafianos.github.io/garment3dgen

### v1

@@ Abstract
We introduce Garment3DGen a new method to synthesize 3D garment assets from a base mesh given a single input image as guidance. Our proposed approach allows users to generate 3D textured clothes based on both real and synthetic images, such as those generated by text prompts. The generated assets can be directly draped and simulated on human bodies. First, we leverage the recent progress of image-to-3D diffusion methods to generate 3D garment geometries. However, since these geometries cannot be utilized directly for downstream tasks, we propose to use them as pseudo ground-truth and set up a mesh deformation optimization procedure that deforms a base template mesh to match the generated 3D target. Second, we introduce carefully designed losses that allow the input base mesh to freely deform towards the desired target, yet preserve mesh quality and topology such that they can be simulated. Finally, a texture estimation module generates high-fidelity texture maps that are globally and locally consistent and faithfully capture the input guidance, allowing us to render the generated 3D assets. With Garment3DGen users can generate the textured 3D garment of their choice without the need of artist intervention. One can provide a textual prompt describing the garment they desire to generate a simulation-ready 3D asset. We present a plethora of quantitative and qualitative comparisons on various assets both real and generated and provide use-cases of how one can generate simulation-ready 3D garments.
Figure 1: We present Garment3DGen, a fully-automated method to transform a base garment mesh to simulation-ready asset directly from images or text prompts. Our method enables rapid asset generation in a frictionless manner, commoditizing content creation which would otherwise require specialized software and expertise. We demonstrate applications such as physics-based simulation and hand-cloth interaction in VR.
@@ 1 Introduction
3D asset creation is the process of designing and generating geometries and materials for 3D experiences. It has direct applications across several industries such as gaming, movies, fashion as well as VR applications. Traditionally, simulation-ready garments are hard to obtain and are created through a laborious time-consuming process requiring specialized software [11, 7, 64] relying on experienced artists. Currently, creating virtual clothing for simulation is a challenging task. Garments need to be manually designed and draped onto an underlying body. Additionally, the topology of the garment needs to take simulation into consideration in order to enable pleasing results. Low-friction asset creation will be the key enabler in the future to unlock virtual applications at scale. Generative AI will be a cornerstone technology that will allow anyone, ranging from novice users to experts, to create customized avatars and to contribute to building personalized virtual experiences. In addition, it will assist in the design process to facilitate faster exploration and creation of new designs.
To tackle this task, we set out to develop a method termed Garment3DGen that creates 3D garments directly from image inputs. Given a base geometry mesh and a single image, Garment3DGen performs topology-preserving mesh-based deformations to match the image guidance and synthesizes new 3D assets on the fly. Our generated garments comprise of posed geometries that stylistically match the input image, and high-resolution texture maps. The provided image guidance can be either from the real world or synthetically generated [56, 12] which enables us to create both real and fantastical 3D garments.
@@
While there are several methods in the literature that have focused on this problem, they all suffer from some key limitations. One way to tackle this problem would be to utilize recent image-to-3D techniques [37]. Given a single image as input, such methods synthesize a specific number of views captured from pre-set viewpoints and then employ multi-view reconstruction techniques to obtain the 3D asset. However, the output geometries tend to be coarse and lack fine-level details due to the use of Marching Cubes to extract the output 3D geometry. Another drawback is that the output garments are watertight, have arbitrary scale making it challenging to drape them om human bodies and simulate them directly as this would require manual intervention to post-process the geometry (e.g., manually create arm/neck/waist holes for a t-shirt geometry as well as re-meshing that would alter the mesh topology). Alternatively, one could follow a NeRF-based approach where a handful of views of the base mesh are utilized to train a NeRF which can be stylized in an iterative manner [24]. Such an approach however does not guarantee multi-view consistency of the newly stylized garment as the NeRF training and stylization are happening in an iterative manner. In addition, it is a time-consuming process and the final geometry is not simulation-ready. Recent works have focused on 3D Gaussian Splatting [29] to generate 3D assets from image inputs. While such methods are fast and of fairly high reconstruction quality, their output splats are hard to be used for any downstream task besides rendering. Another direction of research predicts 2D garment patterns [30, 33, 5, 44] which can be optimized using differentiable simulation. Such approaches generate simulation-ready garments but cannot generalize to fantastical AI-generated garments and cannot benefit from the recent progress of text and image-based diffusion models.
@@
To this end, we carefully designed Garment3DGen to tackle each of the aforementioned challenges: i) reconstruction-based approaches output geometries that are watertight, coarse and the garments cannot be draped on human bodies ii) deformation-based approaches are under-constrained when given a single image/text prompt and their outputs do not faithfully match the provided guidance and iii) simulation-based approaches fail to generalize to new garment types. Garment3DGen is capable of producing high-quality simulation-ready stylized garment assets complete with associated textures. We approach this task from a mesh deformation-based perspective as we believe that it provides better properties and more fine-grained control for the output geometries compared to alternative NeRF-based or reconstruction-based approaches. Mesh-based deformations can preserve the mesh topology which in turn can allow for UV texture transfer, they can preserve the arm/body/head holes of the garment geometry instead of outputting watertight meshes, and can provide output meshes the triangles of which are not distorted and can be draped on human bodies and simulated. Our method takes as inputs a single image and a base template mesh and outputs a deformed mesh that faithfully follows the image guidance while preserving the structure and topology of the base mesh. Our first contribution stems from supervising the mesh deformation process directly in the 3D space which directly allows physics-inspired losses that ensure simulation-readiness instead of solely relying on image-based or embedding-based supervisions [74, 46, 48]. In the absence of 3D ground-truth, we build upon the progress of diffusion-based multi-view consistent image generation to obtain a coarse 3D geometry that can serve as pseudo ground-truth. However, strictly enforcing 3D supervisions using a coarse mesh would result in deformed meshes that lack fine-level details and are not simulation-ready. Thus, we utilize a pre-trained MetaCLIP model [74] since it provides a balanced subset over the metadata distribution and finetune it to garment data while introducing additional supervisions in the image space using differentiable rendering and in the embedding space. Finally, we propose a carefully designed texture estimation module to predict the texture maps required to create the final 3D garment.
@@
We conducted a plethora of experiments that demonstrate that our method can generate 3D garment assets i) directly from images allowing a frictionless experience where users indicate a requested garment by providing a reference image and quickly obtain a corresponding high-quality 3D asset without manual intervention and ii) from textual inputs describing both real and fantastical garments, and iii) even simple garment sketches that one can quickly draw. Moreover, we have developed a body-garment co-optimization framework that enables us to scale and fit the garment to a parametric body model. This allows us to animate the body model and perform physics-based cloth simulation, resulting in a more accurate representation of the garment’s behavior in various novel scenarios. In summary, our contributions are as follows:
@@
  • We propose a new approach for 3D geometry and texture generation for garments given a base mesh and a single image guidance as inputs. We believe that this is the first work aiming to generate textured garment assets that can be useful for downstream simulation tasks.
@@
  • We propose a body-cloth optimization framework that fits the generated 3D garment on a parametric body which unlocks applications such as dressing avatars and the accurate simulation of clothes without artist intervention.
@@ 2 Related work
Garment Modeling: An important line of work is focused on designing[14, 6], capturing [77], registering, reconstructing [54, 78], and representing [66, 39] clothes and their texture [8] from image or video inputs. Cloth registration is an important task as it allows us to fit parametric garment templates to in-the-wild scans or in controlled environments [23] and then use the registered clothes for downstream tasks. [21] learns a shape diffusion-based prior from captured 4D data in order to enable registration of texture-less cloth, whereas [34] aligns the garment geometry to real world captures using a coarse-to-fine method that leverages intrinsic manifold properties with neural deformation fields. Aiming to accurately model clothes and discover new compact ways to represent them, CaPhy [65] recovers a dynamic neural model of clothing in a similar fashion to SNUG [58] by leveraging 3D supervised training in combination with physics-based losses. [73] introduced a physically-inspired appearance representation by learning view-dependent and dynamic shadowing effects. Finally, recent methods have explored modeling clothes using graph neural networks [19, 22, 50]. An alternative way to represent clothes is via sewing patterns as this ensures an efficient representation of developable [62, 57] and manufacturable garment items which can be easily modified. A plethora of works [52, 30, 2, 4] have followed this approach for garment reconstruction [36, 69, 68, 49], generation [59] and draping [32].
@@
Garment Deformation and Stylization: The recent progress in large language and image-to-3D models has unlocked new ways of representing and reconstructing dressed avatars [27, 70] from a single text or image prompt. These methods typically generate a handful of multi-view consistent views [38, 72, 37, 53, 40] given a single image or text input or directly optimize a 3D scene [51] using a 3D scene parameterization, similar to Neural Radiance Fields [47]. However such methods generate coarse, watertight meshes that in the context of garments do not have the required topology and structure to be draped on humans and simulated [63]. Editing and stylizing 3D surfaces has been explored in the context of optimizing directly on the 3D space [61, 35, 28] and more recently using triplanes [16] and text-to-mesh formulations [10, 48]. For example, [46] performs mesh stylization by predicting color and local geometric details that follow a text prompt. Deformation-based approaches [67, 26, 3, 71, 80, 17, 20, 76] can leverage these foundation models to enforce supervision signals for text and image-based stylization [13] and manipulation [18] of 3D meshes. Recently, a line of work [55, 9, 79, 75] applied text-to-image generation models to create textures based on the mesh and given text/image. Extending such techniques to clothes is a complex task as the supervision signals of a single image or text-prompt are insufficient to ensure that the deformed clothes will be simulation-ready.
@@
Figure 2: Overview: Given an input 3D base mesh and a target garment image we first generate 3D pseudo ground-truth using a diffusion-based method and utilize the output geometry as a soft supervision signal during the deformation process. Our 3D generated geometry preserves the topology and structure of the base mesh as depicted by the colors of the sleeves/collar while accurately capturing the geometry of the input image. The texture-estimation module then outputs the corresponding UV texture that along with the geometry comprise our final generated 3D garment.
3 Method
@@
Our approach takes as input a single image $I$ and a base garment template mesh $M_{\text{in}}$ and performs a topology-preserving deformation of the input geometry given the image guidance to obtain the target deformed mesh $M_{\text{def}}=D(I,M_{\text{in}})$ where $D$ is a function represented by a neural network that optimizes over the input mesh. An overview of our method is depicted in Fig. 2.
@@ 3.1 Target Geometry Generation
We propose to leverage the recent progress of single-image-to-3D methods to obtain a coarse geometry of $I$ denoted by $M_{\text{guide}}(I)$ and use it as much stronger supervision both directly in the 3D space as well as in projected 2D space through differentiable rendering. A cross-domain diffusion model [40] is employed which synthesizes RGB and normal images from six views given the input image $I$ captured from the same predefined viewpoint. A multi-view 3D reconstruction algorithm [83] is then utilized that, given the generated views, it outputs a watertight, relatively coarse geometry $M_{\text{guide}}(I)$ along with vertex colors of the garment in the input image. While this mesh cannot serve as the final simulation-ready result due to its poor mesh quality which is due to Marching Cubes [42] or potential inaccuracies of the multi-view generation. Additionally, the fact that it is watertight prevents us from draping the garment on a body (e.g. due to missing armholes). Nonetheless, it provides useful information to serve as a pseudo ground-truth that serves as a reference to deform $M_{\text{in}}$ towards. Hence, we update the optimization function as follows $M_{\text{def}}=D(I,M_{\text{in}},M_{\text{guide}}(I))$. The alternative approach would be to rely on the input image $I$ as the sole supervision for mesh deformation which would result in a severely under-constrained optimization with low-quality output meshes that are uncanny, over-deformed, and they fail to capture the subtle details of the image guidance. For example, starting from a template T-shirt geometry with guidance of the image of an armor as input, one could extract CLIP embeddings for both $I$ and the renders of $M_{\text{def}}$ following a similar approach to TextDeformer [18]. By enforcing supervisions on the embeddings, the goal is obtain an output mesh that would resemble the requested armor. In practice, the supervision from the embedding of a single exemplar image is not strong enough to produce high quality output.
3.2 Mesh Deformer: Topology-Preserving Deformations
Aiming to preserve the structure and topology of the input base mesh while still allowing for image-base stylizations, we propose an approach which deforms $M_{\text{in}}$ instead of generating a new geometry that would be hard to directly use in downstream tasks like reconstruction-based methods would do. Hence, inspired by Neural Jacobian Fields [1], we parameterize $M_{\text{in}}$ using a set of per-triangle Jacobians which define a deformation. Following the same formulation, for simplicity we represent per-triangle Jacobians as matrices $J_{i}\in R^{3x3}$ and solve a Poisson optimization problem to obtain the deformation map $\Phi^{*}$ as the mapping with Jacobian matrices for each triangle that are closest to $J_{i}$. More formally this is represented as:
@@ 3.2 Mesh Deformer: Topology-Preserving Deformations
where $\nabla(\Phi)$ denotes the Jacobian of $\Phi$ at triangle $t_{i}$, with $|t_{i}|$ being the area of the triangle. We optimize the deformation mapping $\Phi$ indirectly by optimizing the matrices $J_{i}$ which define $\Phi^{*}$. These Jacobians are initialized to identity matrices. With the Jacobian representation at hand, we optimize over the triangles of $M_{\text{in}}$ by introducing a several losses each one of which addressing a specific issue.
@@
Regularizations: We introduce several regularizations on the deformed 3D mesh to ensure that it maintains key properties. First, we introduce Laplacian smoothing [15] denoted by $L_{\text{Lap}}$ to redistribute vertex positions based on the average positions of neighboring vertices. This smoothing process helps to reduce irregularities and improves the overall shape of the mesh. To produce simulation-ready meshes we penalize very small surface area for triangles denoted by $L_{\text{triag}}$ as that would result in meshes that are difficult to simulate. We do this by regularizing the edge length and by minimizing the inverse of the squared sum of the triangle areas. Note that there is a trade-off between how much a mesh can freely deform, (e.g. a shirt becoming a medieval armor with spikes), and how much regularization it requires such that the new garment can be placed on a parametric body and simulated.
2D Supervisions: We utilize a rasterization-based differentiable renderer [31] denoted by $R$ and pass both the deformed mesh $M_{\text{def}}$ in each iteration and target pseudo ground-truth mesh $M_{\text{guide}}(I)$ to obtain $K$ image renders:
${I_{\text{def}}}_{i}=R(M_{\text{def}},C_{i}),\;\;\;i=1\dots K,$ (3)
from randomly sampled camera views $C_{i}$. With ${I_{\text{I}}}_{i}$ computed in a similar fashion for $M_{\text{guide}}(I)$ we employ the L1 loss between the deformed and target renders:
@@
Embedding Supervisions: We observe that passing garment images through a pre-trained MetaCLIP model [74] results in deformed output garments that are overly distorted, uncanny and fail to capture the fine-level details provided in the input images. This is due to the weak supervision signal contained in these embeddings. Likewise, this holds true for other mesh classes that are not well represented in the data on which MetaCLIP was trained on, such as humans. This is because MetaCLIP fails to capture the subtle differences between different garments, their properties, and materials. To overcome this limitation, we propose to use garment-specific embeddings obtained fine-tuned on fashion data termed as FashionCLIP. The latent space for this model is better tuned for fashion concepts and as a result, the embeddings provide a stronger guidance for the deformations. We represent this loss as:
$L_{\text{E}}=\frac{1}{K}\sum_{i=1}^{K}\text{CosSim}\left(\text{FashionCLIP}\big{(}{I_{\text{def}}}_{i}\big{)},\text{FashionCLIP}\big{(}{I_{\text{I}}}_{i}\big{)}\right),$ (5)
where CosSim is the cosine similarity. This embedding loss acts as a soft supervision signal between the embeddings of the deformed mesh $M_{\text{def}}$ and those of the pseudo-ground-truth $M_{\text{guide}}(I)$. This behavior is desired since we aim to benefit from the embedding representations of the MetaCLIP model and sufficiently deform the base mesh towards the target without capturing all of its shortcomings. For example we need to preserve the arm/head holes of the base geometry while the target mesh is watertight, or capture the fine-level details depicted in the input image that $M_{\text{guide}}(I)$ might have failed to represent well. In summary, the total loss $L_{\text{T}}$ is defined as follows where $w_{*}$ is the corresponding weight for each loss:
@@
where the corresponding weights aim to strike a balance between utilizing the 3D pseudo ground-truth with the CD loss while capturing the finer image details through the embedding and 2D losses. We provide additional details in the supplementary material.

### v2

# Garment3DGen: 3D Garment Stylization and Texture Generation

@@ Abstract
We introduce Garment3DGen a new method to synthesize 3D garment assets from a base mesh given a single input image as guidance. Our proposed approach allows users to generate 3D textured clothes based on both real and synthetic images, such as those generated by text prompts. The generated assets can be directly draped and simulated on human bodies. We leverage the recent progress of image-to-3D diffusion methods to generate 3D garment geometries. However, since these geometries cannot be utilized directly for downstream tasks, we propose to use them as pseudo ground-truth and set up a mesh deformation optimization procedure that deforms a base template mesh to match the generated 3D target. Carefully designed losses allow the base mesh to freely deform towards the desired target, yet preserve mesh quality and topology such that they can be simulated. Finally, we generate high-fidelity texture maps that are globally and locally consistent and faithfully capture the input guidance, allowing us to render the generated 3D assets. With Garment3DGen users can generate the simulation-ready 3D garment of their choice without the need of artist intervention. We present a plethora of quantitative and qualitative comparisons on various assets and demonstrate that Garment3DGen unlocks key applications ranging from sketch-to-simulated garments or interacting with the garments in VR. Code is publicly available.
Figure 1: Garment3DGen, automatically transforms a base garment mesh to simulation-ready asset directly from images or text in a frictionless manner unlocking applications such as cloth and hand-cloth interaction in a VR.
@@ 1 Introduction
3D asset creation is the process of designing and generating geometries and materials for 3D experiences. It has direct applications across several industries such as gaming, movies, fashion as well as VR applications. Traditionally, simulation-ready garments are hard to obtain and are created through a laborious time-consuming process requiring specialized software [14, 8, 76] relying on experienced artists. Currently, creating virtual clothing for simulation is a challenging task. Garments need to be manually designed and draped onto an underlying body. Additionally, the topology of the garment needs to take simulation requirements into consideration in order to enable stable and visually pleasing results. Low-friction asset creation will be the key enabler to unlock virtual applications at scale. Generative AI will be a cornerstone technology that will allow anyone, from novice users to experts, to create customized avatars and to contribute to building personalized virtual experiences. In addition, it will assist in the design process to facilitate faster exploration and creation of new designs.
@@
To tackle this task, we set out to develop a method that creates 3D garments directly from images. Given a base geometry mesh and a single image, Garment3DGen performs topology-preserving mesh-based deformations to match the image guidance and synthesizes new 3D assets on the fly. Our generated garments comprise of posed geometries that stylistically match the input image, and high-resolution texture maps. The provided image guidance can be either from the real world or synthetically generated [66, 16] which enables us to create both real and fantastical 3D garments. One way to tackle this problem would be to utilize recent image-to-3D techniques [47]. Given a single image as input, such methods synthesize a specific number of views captured from pre-set viewpoints and then employ multi-view reconstruction techniques to obtain the 3D asset. However, the output geometries tend to be coarse and lack fine-level details due to the use of Marching Cubes to extract the output 3D geometry. Another limitation is that the output garments are watertight and have arbitrary scale, making it difficult to drape them on human bodies and simulate them directly. This is because manual intervention would be required to post-process the geometry. For example, creating arm, neck, and waist holes for a t-shirt geometry, as well as re-meshing to produce mesh topologies suitable for downstream applications which are sensitive to poor mesh qualities. Alternatively, one could follow a NeRF-based approach where a handful of views of the base mesh are utilized to train a NeRF which can be stylized in an iterative manner [30]. However, such an approach does not guarantee multi-view consistency of the newly stylized garment because the NeRF training and stylization are happening in an iterative manner. In addition, it is a time-consuming process and the final result is not suitable for simulation. Recent works have focused on 3D Gaussian Splatting [36] to generate 3D assets from image inputs. While such methods are fast and of produce high reconstruction quality, their output splats are hard to be used for any downstream task besides rendering. Another direction of research predicts 2D garment patterns [37, 5, 17] which can be optimized using differentiable simulation [43, 11]. Such approaches generate simulation-ready garments but cannot generalize to fantastical AI-generated garments and are constrained to specific garment styles. For example, when testing SewFormer [46] on upper-body garments, it only works well on input with V-neck shirts, which is very limiting; while for other garments (long-sleeve with round neck) it fails to generate accurate patterns.
@@
To this end, we carefully designed Garment3DGen to tackle each of the aforementioned challenges: i) reconstruction-based approaches output geometries that are watertight, coarse and the garments cannot be draped on human bodies ii) deformation-based approaches are under-constrained when given a single image or text prompt and their outputs do not faithfully match the provided guidance and iii) simulation-based approaches fail to generalize to new garment types. Garment3DGen produces high-quality simulation-ready stylized garment assets along with the associated textures. We approach this task from deformation-based perspective as we believe that it provides better properties and more fine-grained control for the output geometries compared to alternative NeRF-based or reconstruction-based approaches. Mesh-based deformations can preserve the mesh topology which in turn can allow for UV texture transfer, they can preserve the arm/body/head holes of the garment geometry instead of outputting watertight meshes, and can provide output meshes the triangles of which are not distorted and can be draped on human bodies and simulated. Our method takes as inputs a single image and a base template mesh and outputs a deformed mesh that faithfully follows the image guidance while preserving the structure and topology of the base mesh. Our first contribution stems from supervising the mesh deformation process directly in the 3D space which directly allows physics-inspired losses that ensure simulation-readiness instead of solely relying on image-based or embedding-based supervisions [87, 56, 58]. In the absence of 3D ground-truth, we build upon the progress of diffusion-based multi-view consistent image generation to obtain a coarse 3D geometry that can serve as pseudo ground-truth. However, strictly enforcing 3D supervisions using a coarse mesh results in deformed meshes that lack fine-level details and are not simulation-ready. Thus, we utilize a pre-trained MetaCLIP model [87] finetuned to garment data and introduce additional losses in the image as well as the embedding spaces using differentiable rendering. Finally, we propose a carefully designed texture estimation module to predict the texture maps required to create the final 3D garment.
@@
We conducted a plethora of experiments that demonstrate that our method generates 3D garment assets i) directly from images allowing a frictionless experience where users indicate a requested garment by providing a reference image and quickly obtain a high-quality 3D asset without manual intervention and ii) from text describing both real and fantastical garments, and iii) even garment sketches that one can quickly draw. Moreover, we have developed a body-garment co-optimization framework that enables us to scale and fit the garment to a parametric body model. This allows us to animate the body model and perform physics-based cloth simulation, resulting in a more accurate representation of the garment’s behavior in various novel scenarios. In summary, our contributions are as follows:
@@
  • We propose a new deformation-based approach for 3D geometry and texture generation for garments given a base mesh and a single image guidance as inputs.
@@ 2 Related Work
Garment Modeling: An important line of work is focused on designing [19, 7, 83], capturing [92], registering [28], reconstructing [64, 15, 59, 34, 89, 42], and representing [78, 50] clothes and their texture [9] from images or videos. BCNet [34] predicts garment vertex displacements whereas Moon et al. [59] map 3D garments to a parametric body using dense keypoints. Both these approaches cannot generalize to loose clothes moving freely. Guo et al. [27] learns a shape diffusion-based prior from captured 4D data in order to enable registration of texture-less cloth, whereas Lin et al. [44] aligns the garment geometry to real world captures using a coarse-to-fine method that leverages intrinsic manifold properties with neural deformation fields. Aiming to model clothes and discover compact ways to represent them, CaPhy [77] recovers a dynamic neural model of clothing similar to SNUG [68] by leveraging 3D supervised training in combination with physics-based losses. Xiang et al. [86] introduced a physically-inspired appearance representation by learning view-dependent and dynamic shadowing effects. Finally, recent methods explored modeling clothes using graph neural networks [25, 29, 60]. An alternative way to represent clothes is via sewing patterns as it ensures an efficient representation of developable [74, 67] and manufacturable garments which can be easily modified. A plethora of works [62, 2, 4] have followed this path for garment reconstruction [46, 81, 80], generation [70, 31, 49] and draping [41].
@@
Garment Deformation and Stylization: The recent progress in language and image-to-3D models has unlocked new ways of representing and reconstructing dressed avatars [35, 82] from text or images. Such methods typically generate multi-view consistent views [48, 85, 47, 63, 51] given a single image or text input or directly optimize a 3D scene [61] using a 3D scene parameterization, similar to Neural Radiance Fields [57]. However such methods generate coarse, watertight meshes that in the context of garments do not have the required topology and structure to be draped on humans and simulated [75]. Editing and stylizing 3D surfaces has been explored for optimizing directly in 3D [72, 45] and more recently using triplanes [21] and text-to-mesh formulations [12, 58]. For example, Michel et al. [56] performs mesh stylization by predicting color and local geometric details that follow a text prompt. Deformation-based approaches [54, 79, 33, 3, 84, 94, 23, 26, 91, 22] can leverage these foundation models to enforce supervision for text and image-based stylization [18] and manipulation [24] of 3D meshes. Recent work [65, 10, 93, 90] applied text-to-image generation models to create textures based on the mesh and given text/image. Extending such techniques to clothes is a complex task as the supervision signals of a single image or text-prompt are insufficient to ensure that the deformed clothes will be simulation-ready.
@@
Figure 2: Overview: Given an input 3D base mesh and a target garment image we first generate 3D pseudo ground-truth using a diffusion-based method and utilize the output geometry as a soft supervision signal during the deformation process. Our 3D generated geometry preserves the topology and structure of the base mesh as depicted by the colors of the sleeves/collar while accurately reflecting the geometry and details of the input image. Finally, we introduce a texture-estimation module which outputs the corresponding UV texture that along with the geometry comprise our final generated 3D garment.
3 Methodology
@@
Our approach takes as input a single image $I$ and a base garment template mesh $M_{\text{in}}$ and performs a topology-preserving deformation of the input geometry given the image guidance to obtain the target deformed mesh $M_{\text{def}}=D(I,M_{\text{in}})$ where $D$ is a function that optimizes over the input mesh. An overview of our method is depicted in Fig. 2.
@@ 3.1 Target Geometry Generation
We propose to leverage the recent progress of single-image-to-3D methods to obtain a coarse geometry of $I$ denoted by $M_{\text{guide}}(I)$ and use it as much stronger supervision both directly in the 3D space as well as in projected 2D space through differentiable rendering. A cross-domain diffusion model [71, 51] is employed which synthesizes RGB and normal images from six views given the input image $I$ captured from the same predefined viewpoint. A multi-view 3D reconstruction algorithm based on the LRM architecture [88] is then utilized that, given the generated views, it outputs a watertight, relatively coarse geometry $M_{\text{guide}}(I)$ of the garment in the input image. Such meshes cannot serve as the final simulation-ready result due to its poor mesh quality which is due to Marching Cubes [53] (or FlexiCubes [69]) or potential inaccuracies of the multi-view generation. Additionally, the fact that it is watertight prevents us from draping the garment on a body (e.g. missing armholes). Nonetheless, it provides useful information to serve as a pseudo ground-truth that serves as a reference to deform $M_{\text{in}}$ towards. Hence, we update the optimization function as follows $M_{\text{def}}=D(I,M_{\text{in}},M_{\text{guide}}(I))$. The alternative approach would be to rely on the input image $I$ as the sole supervision for mesh deformation which would result in a severely under-constrained optimization with low-quality output meshes that are uncanny, over-deformed, and they fail to capture the subtle details of the image guidance. For example, starting from a template T-shirt geometry with guidance of the image of an armor, one could extract CLIP embeddings for both $I$ and the renders of $M_{\text{def}}$ following a similar approach to TextDeformer [24]. By enforcing supervisions on the embeddings, the goal is obtain an output mesh that would resemble the requested armor. In practice, the supervision from the embedding of a single exemplar image is not strong enough to produce high quality output.
3.2 Topology-Preserving Deformations
In order to preserve the structure and topology of the input base mesh while enabling image-base stylizations, we propose an approach which deforms $M_{\text{in}}$. In contrast, generating novel geometry using reconstruction-based methods would be hard to use directly in downstream tasks. Hence, inspired by Neural Jacobian Fields [1], we parameterize $M_{\text{in}}$ using a set of per-triangle Jacobians which define a deformation. Following the same formulation, for simplicity we represent per-triangle Jacobians as matrices $J_{i}\in R^{3x3}$ and solve a Poisson optimization problem to obtain the deformation map $\Phi^{*}$ as the mapping with Jacobian matrices for each triangle that are closest to $J_{i}$. More formally this is represented as:
@@ 3.2 Topology-Preserving Deformations
where $\nabla(\Phi)$ denotes the Jacobian of $\Phi$ at triangle $t_{i}$, with $|t_{i}|$ being the area of the triangle. We optimize the deformation mapping $\Phi$ indirectly by optimizing the matrices $J_{i}$ which define $\Phi^{*}$. These Jacobians are initialized to identity matrices. With the Jacobian representation at hand, we optimize over the triangles of $M_{\text{in}}$ by introducing a several losses each addressing a specific issue.
@@
Regularizations: We introduce several regularizations on the deformed 3D mesh to ensure that it maintains key properties. First, we introduce Laplacian smoothing [20] denoted by $L_{\text{Lap}}$ to redistribute vertex positions based on the average positions of neighboring vertices. This smoothing process helps to reduce irregularities and improves the overall mesh shape. To produce simulation-ready meshes we penalize very small surface area triangles denoted by $L_{\text{triag}}$ as that would result in meshes that are difficult to simulate. We do this by regularizing the edge length and by minimizing the inverse of the squared sum of the triangle areas. Note that there is a trade-off between how much a mesh can freely deform, (e.g. a shirt becoming a spiky medieval armor), and how much regularization it requires such that the garment can be placed on a parametric body and simulated.
2D Supervisions: We utilize a rasterization-based differentiable renderer [38] denoted by $R$ and pass both the deformed mesh $M_{\text{def}}$ in each iteration and target pseudo ground-truth mesh $M_{\text{guide}}(I)$ to obtain $K$ image renders ${I_{\text{def}}}_{i}=R(M_{\text{def}},C_{i}),\;\;\;i=1\dots K$ from randomly sampled camera views $C_{i}$. With ${I_{\text{I}}}_{i}$ computed in a similar fashion for $M_{\text{guide}}(I)$ we employ the L1 loss between the deformed and target renders:
@@
Embedding Supervisions: We observe that passing garment images through a pre-trained CLIP model results in deformed output garments that are overly distorted, uncanny and fail to capture the fine-level details provided in the input images. This is due to the weak supervision signal contained in these embeddings. Likewise, this holds true for other mesh classes that are not well represented in the data on which CLIP was trained on, such as human geometries for example. This is because CLIP fails to capture the subtle differences between garments, their properties and materials. To overcome this limitation, we propose to use garment-specific embeddings obtained from a CLIP model fine-tuned on fashion data named FashionCLIP [13] denoted by FCLIP. The latent space for this model is better tuned for fashion concepts and as a result, the embeddings provide a stronger guidance for the deformations. We represent this loss as:
$L_{\text{E}}=\frac{1}{K}\sum_{i=1}^{K}\text{Cos}\left(\text{FCLIP}\big{(}{I_{\text{def}}}_{i}\big{)},\text{FCLIP}\big{(}{I_{\text{I}}}_{i}\big{)}\right),$ (4)
where CosSim is the cosine similarity. This embedding loss acts as a soft supervision signal between the embeddings of the deformed mesh $M_{\text{def}}$ and those of the pseudo-ground-truth $M_{\text{guide}}(I)$. This behavior is desired since we aim to benefit from the embedding representations of the CLIP model and sufficiently deform the base mesh towards the target without capturing all of its shortcomings. For example we need to preserve the arm/head holes of the base geometry while the target mesh is watertight, or capture the fine-level details depicted in the input image that $M_{\text{guide}}(I)$ might have failed to represent well. In summary, the total loss $L_{\text{T}}$ is defined as follows where $w_{*}$ is the corresponding weight for each loss:
@@
where the corresponding weights balance between utilizing the 3D pseudo ground-truth with the CD loss while capturing the finer details through the embedding and 2D losses.